OpenStudy (anonymous):

Determine the sum or difference in clock 12 arithmetic. (7 + 10) + (8 + 7)

OpenStudy (anonymous):

ten hours after seven o'clock is 5 right?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Yes...

OpenStudy (anonymous):

actually if you want to do this the easy way, add all the numbers and then subtract 12 repeated until you get a number between one and 12

OpenStudy (anonymous):

So the answer is 8?

OpenStudy (anonymous):

\[7+10+8+7=32\] subtract 12 twice and get 8 yes

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

or you could say \[7+10=5\] \[5+8=1\] \[1+7=8\]